OpenStudy (anonymous):

In a word tile game, you score one positive point for each letter tile that you use and one negative point for each tile that you have left. 10. During one round, Shelley used 14 tiles and could not use 6. What was her score for that round? 11. Pete used 4 tiles, but he could not use 8. What was his score that round? 12. Reasoning In the game, if you have 18 tiles and you cannot use 3 of them, what will your score be for that round? Explain how you found the answer.

OpenStudy (anonymous):

10) 14-6=8 11) 4-8=-4 12)18-3=15